A chemist mixes a 8% alcohol solution with a 14% alcohol solution to make 600 milliliters of a 12% alcohol solution. How many milliliters of each solution does the chemist use?
I need help on how to set the problem up and the answer. If possible, can you show your work, step by step?

Answer by ewatrrr(24785)


Hi
Let x and (600-x)represent the amounts of the 14% and 8% solutions respectively
Question states***
.14x + .08(600-x) = .12*600ml
Solving for x
.06x = .04*600ml x = 400ml, amount of 14%, 200ml of 8%. (600ml-400ml))
CHECKING our Answer***
.14*400 + .08*200 = 56 + 16 = 72 = .12*60